OpenStudy (anonymous):

please help..schools almost over. 3/x-5 + 5/x^2 -25

jimthompson5910 (jim_thompson5910):

Can you factor x^2 -25 at all?

OpenStudy (anonymous):

x= -5 x= 5?

jimthompson5910 (jim_thompson5910):

You're close, it's really (x-5)(x+5)

OpenStudy (anonymous):

Ok.. i still dont know the anwser :[

jimthompson5910 (jim_thompson5910):

In order to add the two fractions, what must the denominators be?

jimthompson5910 (jim_thompson5910):

Any ideas?

OpenStudy (anonymous):

no

jimthompson5910 (jim_thompson5910):

Alright, here's how you solve this problem

OpenStudy (anonymous):

k thank you!

jimthompson5910 (jim_thompson5910):

\[\Large \frac{3}{x-5}+\frac{5}{x^{2}-25} \] \[\Large \frac{3}{x-5}+\frac{5}{\left(x-5\right)\left(x+5\right)} \] \[\Large \frac{3\left(x+5\right)}{\left(x-5\right)\left(x+5\right)}+\frac{5}{\left(x-5\right)\left(x+5\right)} \] \[\Large \frac{3x+15}{\left(x-5\right)\left(x+5\right)}+\frac{5}{\left(x-5\right)\left(x+5\right)} \] \[\Large \frac{3x+15+5}{\left(x-5\right)\left(x+5\right)} \] \[\Large \frac{3x+20}{\left(x-5\right)\left(x+5\right)} \] \[\Large \frac{3x+20}{x^{2}-25} \] So \[\Large \frac{3}{x-5}+\frac{5}{x^{2}-25} \] simplifies to \[\Large \frac{3x+20}{x^{2}-25} \]

jimthompson5910 (jim_thompson5910):

I'm basically doing the following Step 1) Get all the denominators equal Step 2) Combine the numerators over the common denominator

OpenStudy (anonymous):

then the anwser is 3x+20/x-5 x+5?

jimthompson5910 (jim_thompson5910):

It's either \[\Large \frac{3x+20}{\left(x-5\right)\left(x+5\right)} \] or it is \[\Large \frac{3x+20}{x^{2}-25} \] Either one works since the two are equivalent. The book might like one answer over the other though.

jimthompson5910 (jim_thompson5910):

Does that make sense?

OpenStudy (anonymous):

Yup

jimthompson5910 (jim_thompson5910):

that's great